# Structural adaptation via directional regularity: rate accelerated estimation in multivariate functional data

### v2

@@ 6 Discussion
In this section we discuss some key aspects of our methodology. Extensions of our approach to more general settings, such as the random design framework with heteroscedastic noise, can be found in the Supplementary Material.
@@ 6.1 Computational Considerations
The computational complexity of Algorithm 1 is driven primarily by the identification process, due to the additional loop over the grid of spacings $\mathbf{\Delta}$. Since we are performing a linear grid search in each basis vector containing $\sqrt{M_{0}}$ points, the complexity of nearest neighbor interpolation for each surface is $O(M_{0})$, resulting in $O(N\times M_{0})$ for $N$ sample paths. By searching over $M_{0}$ points to determine the closest point to each $\mathbf{t}_{m}$ in terms of $\ell_{1}$ norm, the complexity of computing $\widehat{\sigma}^{2}$ is $O(N\times M_{0})$, so the complexity of $\widehat{H}_{\mathbf{v}(\beta)}(\mathbf{t},\Delta)$ is $O(N\times M_{0})$ for each $\mathbf{t}$, $\Delta$. The complexity of the full identification process is therefore $O(N\times\#\widetilde{\mathcal{T}}\times K_{0}\times M_{0})$. In the context of fda, a complexity of at least $O(N\times\#\widetilde{\mathcal{T}}\times M_{0})$ is to be expected. Since $K_{0}$ is relatively small (e.g 15 points), our algorithm is reasonably good in terms of computational complexity, considering the possible gains in the rates of convergence.
@@ 7.1 Smoothing Bivariate Functional Data
Observe that Fubini’s theorem implies $\mathcal{R}\left(\textbf{B},M_{0}\right)=\|\mathcal{R}\left(\cdot,\textbf{B},M_{0}\right)\|,$ so the integrated risk can readily be recovered from the pointwise risk. Proposition 5 provides a pointwise risk bound. Assumptions can be found in the Appendix.

### v3

@@ 6 Discussion
In this section we discuss some key aspects of our methodology.
@@ 6.1 Computational Considerations
The computational complexity of Algorithm 1 is driven primarily by the identification process, due to the additional loop over the grid of spacings $\boldsymbol{\Delta}$. Since we are performing a linear grid search in each basis vector containing $\sqrt{M_{0}}$ points, the complexity of nearest neighbor interpolation for each surface is $O(M_{0})$, resulting in $O(N\times M_{0})$ for $N$ sample paths. By searching over $M_{0}$ points to determine the closest point to each $\mathbf{t}_{m}$ in terms of $\ell_{1}$ norm, the complexity of computing $\widehat{\sigma}^{2}$ is $O(N\times M_{0})$, so the complexity of $\widehat{H}_{\mathbf{v}(\beta)}(\mathbf{t},\Delta)$ is $O(N\times M_{0})$ for each $\mathbf{t}$, $\Delta$. The complexity of the full identification process is therefore $O(N\times\#\widetilde{\mathcal{T}}\times K_{0}\times M_{0})$. In the context of fda, a complexity of at least $O(N\times\#\widetilde{\mathcal{T}}\times M_{0})$ is to be expected. Since $K_{0}$ is relatively small (e.g 15 points), our algorithm is reasonably good in terms of computational complexity, considering the possible gains in the rates of convergence.
@@ 6.2 Singularity along the maximizing direction
6.3 Random design
In the case of random design, where the surfaces are observed at random, discrete points, our methodology can be applied by first smoothing the surfaces. Only mild moment conditions are required on the rate of approximations.
@@ 7.1 Smoothing Bivariate Functional Data
Observe that Fubini’s theorem implies $\mathcal{R}\left(\textbf{B},M_{0}\right)=\|\mathcal{R}\left(\cdot,\textbf{B},M_{0}\right)\|_{1},$ so the integrated risk can readily be recovered from the pointwise risk. Proposition 5 provides a pointwise risk bound. Assumptions can be found in the Appendix.